# J1: Incentivizing Thinking in LLM-as-a-Judge via Reinforcement Learning

### v1

@@ 5 Related Work
Reward Models. Reward models have been instrumental in both training-time (Ouyang et al., 2022; Lambert et al., 2024) and test-time (Snell et al., 2025) alignment of Large Language Models. Traditional reward models are typically trained with the Bradley-Terry objective and output a scalar score indicating the reward of the response. This design frequently results in poor calibration and generalization across different prompts and responses (Sun et al., 2025; Zhang et al., 2025). Furthermore, such discriminative models do not fully leverage the generative capabilities of LLMs and therefore cannot be scaled up at test time, e.g., with long chain-of-thought or multiple generations (Wang et al., 2024b; Shiwen et al., 2024). As a potential solution, generative reward models have emerged, which we discuss below.
@@
LLM-as-a-Judge and Generative Reward Models. LLM-as-a-Judge and Generative Reward Models (GRMs) conceptually share a similar motivation – the language modeling head in LLMs is used to also output chain-of-thought (CoT) reasoning (in the form of critiques) before generating preference judgments or rewards (Kim et al., 2024a, b; Ankner et al., 2024; Mahan et al., 2024; Ye et al., 2024; Yu et al., 2025; Zhang et al., 2025; Saha et al., 2025). Rewards in such models could either come from training a separate reward head (typically done in GRMs) or from the LM head itself generating real-valued scores as tokens (typically done in LLM-as-a-Judge). Prior work building LLM judges has mostly relied on either prompting (Zheng et al., 2023; Saha et al., 2024), rejection finetuning on self-generated CoTs (Wang et al., 2024d), or preference finetuning on CoT pairs using DPO (Mahan et al., 2024; Trivedi et al., 2024; Saha et al., 2025).
@@
Recently, in some concurrent studies, DeepSeek-GRM (Liu et al., 2025b), JudgeLRM (Chen et al., 2025a), and RM-R1 (Chen et al., 2025b) use Reinforcement Learning for building reasoning judge models. We compare J1 to these methods and show that J1 achieves superior performance with 10$\times$ less data. This is achieved via J1’s methodical novelty that span several axes. First, it is a training recipe that is based only on synthetic data. Second, it focuses on mitigating position bias (a long-standing issue in LLM-as-a-Judge development), leading to the proposal of novel consistency rewards and Pointwise-J1 models trained with pairwise supervision only. Consequently, we are able to comprehensively study different J1 variants that vary across sizes, modeling choices, seed prompts, and reward strategies, enabling us to draw important conclusions on building generalist thinking-judge models with state-of-the-art results.
@@
Reinforcement Learning with Verifiable Rewards. J1 draws inspiration from the recent advancements in improving reasoning through Reinforcement Learning with verifiable rewards. Online optimization algorithms such as GRPO, when combined with accurate and robust rewards, have been shown to elicit enhanced reasoning in LLMs (Guo et al., 2025; Team et al., 2025). In our approach, we construct preference pairs and assign verifiable rewards based on the correctness of the model’s judgments. By optimizing over the generated thinking traces, J1 encourages LLMs to spend more test-time compute before deriving scores and judgments.
@@ 6 Conclusion
We proposed J1, an RL recipe for training Thinking-LLM-as-a-Judge models. Our key innovation was in converting the judgment task into a verifiable task for all kinds of task prompts, themselves both verifiable and non-verifiable, and then optimizing the thoughts and judgments using an online RL method. We trained J1-Llama-8B and J1-Llama-70B, two generalist judge models that outperformed all baselines at their respective model sizes, o1-mini, and even a much larger R1 model on non-verifiable tasks. Using only pairwise supervision, we also trained Pointwise-J1 models that proved to be effective in mitigating position bias, thereby highlighting the potential of both Pairwise and Pointwise Thinking-LLM-as-a-Judge.
Appendix A Prompt Templates
@@
Appendix B Experimental Setup
@@
For training, the policy actor generates 5 rollouts per prompt using ancestral sampling with temperature $1.0$. Training regime uses a learning rate of $61e-$ 6 (decayed to $73e-$ 7 in later steps for pairwise J1-Llama-70B), and a total batch size of 512. The maximum sequence length is set to 4096 tokens for both inputs and outputs.
@@
We experimented with different KL coefficients from $\{0.1,0.01,0.001,0\}$ for J1-Llama-8B, and selected $0.01$ as the best-performing value based on development set accuracy. For J1-Llama-70B, we set the KL coefficient to 0 to encourage more exploration. Preliminarily experiment with entropy bonus during training showed degraded performance: the model tends to generate longer but more repetitive output. See Table 6 in Appendix C for comparison of KL penalty and entropy bonus.
We use 8$\times$×A100 and 64$\times$×A100 GPUs to train J1-Llama-8B, and J1-Llama-70B respectively, and 8$\times$×A100 for inference. Tensor parallelism is set to 8 for both training and inference. During inference, we maintain a maximum generation length of 4096 tokens. The default generation uses greedy decoding, and for inference-time scaling we use sampling with top-p of $0.95$ and temperature of $1$.
Appendix C Additional Results
@@
Comparison of Different Pairwise-J1 Models. In Table 5, we compare the three Pairwise-J1 variants, that as part of the final answer, generate either: (i) only the final verdict, (ii) only real-valued scores for both responses, or (iii) both. We observe that predicting the verdict (without the scores) performs as well as other variants. Having access to scores, however, has other advantages e.g., in quantifying the degree of preference or to rank across multiple responses.
Pairwise-J1 8B Variants PPE Preference PPE Correctness RewardBench JudgeBench RM-Bench FollowBenchEval
w/ Verdict : $(x,a,b)\rightarrow(t,y)$ 60.3 59.2 85.7 42.0 73.4 48.3
w/ Scores: $(x,a,b)\rightarrow(t,s_{a},s_{b})$ 60.7 59.7 85.8 41.7 72.3 46.3
w/ Scores&Verdict: $(x,a,b)\rightarrow(t,s_{a},s_{b},y)$ 59.7 59.3 85.1 41.4 71.5 41.0
Pairwise-J1 8B Variants PPE Preference PPE Correctness RewardBench JudgeBench RM-Bench FollowBenchEval
w/ KL Penalty 60.3 59.2 85.7 42.0 73.4 48.3
w/o KL Penalty 59.8 59.4 84.9 42.9 69.8 48.3
w/o Entropy Bonus 60.3 59.2 85.7 42.0 73.4 48.3
w/ Entropy Bonus 59.3 58.9 84.1 39.4 71.7 42.4
Pointwise-J1 Variants PPE PPE PPE Correctness
Overall Preference Overall MMLU-Pro MATH GPQA MBPP-Plus IFEval
J1-Llama-8B - Greedy 54.5 54.9 53.8 57.0 63.2 52.1 49.8 46.9
J1-Llama-8B - Sampling 55.8 55.7 55.8 58.4 66.5 52.4 52.4 49.8
J1-Llama-8B (Mean-Score@32) 60.4 60.1 60.7 62.0 71.6 56.6 60.6 52.8
J1-Llama-70B - Greedy 57.8 50.6 65.0 73.4 77.4 58.9 60.7 54.6
J1-Llama-70B - Sampling 57.9 50.8 64.9 74.0 79.7 58.2 55.7 59.7
J1-Llama-70B (Mean-Score@32) 70.0 65.1 74.8 81.2 87.6 67.3 81.9 70.8
Table 7: Test-time scaling of Pointwise-J1 models. Judgments are made based on the average scores of the responses.
Models #Pref. Pairs Overall Chat Chat-Hard Safety Reasoning
@@
Skywork-Critic-Llama-3.1-70B (Shiwen et al., 2024) 80K 93.3 96.6 87.9 93.1 95.5
DeepSeek-GRM-27B (Liu et al., 2025b) 237K 86.0 94.1 78.3 88.0 83.8
DeepSeek-GRM-27B (MetaRM voting@32) (Liu et al., 2025b) 237K 90.4 95.3 85.7 89.5 91.0
EvalPlanner-Llama-8B 22K 83.0 85.5 84.0 83.4 79.3
EvalPlanner-Llama-70B 22K 93.8 97.7 89.5 91.7 96.1
J1-Llama-8B 22K 85.7 92.9 80.3 85.6 83.9
J1-Llama-70B 22K 93.3 96.1 90.1 91.9 94.9
Table 8: Results on RewardBench comparing our Pairwise-J1 models with other state-of-the-art models. $\dagger$: Results taken from either RewardBench leaderboard or the corresponding paper.

### v2

@@ 5 Related Work
Reward Models. Reward Models have been instrumental in both training-time (Ouyang et al., 2022; Lambert et al., 2025) and test-time (Snell et al., 2025) alignment of LLMs. Traditional reward models are typically trained with the Bradley-Terry objective and output a scalar score indicating the reward of the response. This design frequently results in poor calibration and generalization across different prompts and responses (Sun et al., 2025; Zhang et al., 2025). Furthermore, such discriminative models do not fully leverage the generative capabilities of LLMs and therefore cannot be scaled up at test time, e.g., with long chain-of-thought or multiple generations (Wang et al., 2024b; Shiwen et al., 2024). As a potential solution, generative reward models have emerged, which we discuss below.
@@
LLM-as-a-Judge and Generative Reward Models. LLM-as-a-Judge and Generative Reward Models (GRMs) conceptually share a similar motivation – the language modeling head in LLMs is used to also output chain-of-thought (CoT) reasoning (in the form of critiques) before generating preference judgments or rewards (Kim et al., 2024a; b; Ankner et al., 2024; Mahan et al., 2024; Ye et al., 2025; Yu et al., 2025b; Zhang et al., 2025; Saha et al., 2025). Rewards in such models could either come from training a separate reward head (typically done in GRMs) or from the LM head itself generating real-valued scores as tokens (typically done in LLM-as-a-Judge). Prior work building LLM judges has mostly relied on either prompting (Zheng et al., 2023; Saha et al., 2024), rejection finetuning on self-generated CoTs (Wang et al., 2024d), or preference finetuning on CoT pairs using DPO (Mahan et al., 2024; Trivedi et al., 2024; Saha et al., 2025; Yu et al., 2025a).
@@
Recently, in some concurrent studies, methods like DeepSeek-GRM (Liu et al., 2025b), JudgeLRM (Chen et al., 2025a), RM-R1 (Chen et al., 2025b), and Reward Reasoning Model (Guo et al., 2025b) use Reinforcement Learning for building reasoning judge models. We compare J1 to these methods (in Tables 1 and Table 4) and show that J1 achieves superior performance with significantly less data. This is achieved via J1’s methodical novelty that span several axes. First, it is a training recipe that is based only on synthetic data. Second, it focuses on mitigating position bias (a long-standing issue in LLM-as-a-Judge development) and multitask learning for a single generalist judge, leading to the proposal of novel consistency rewards and Pointwise-J1 models trained with pairwise supervision only. Consequently, we are able to comprehensively study different J1 variants that vary across sizes, modeling choices, seed prompts, and reward strategies, enabling us to draw important conclusions on building generalist thinking-judge models with state-of-the-art results.
@@
Reinforcement Learning with Verifiable Rewards. J1 draws inspiration from the recent advancements in improving reasoning through Reinforcement Learning with verifiable rewards. Online optimization algorithms such as GRPO, when combined with accurate and robust rewards, have been shown to elicit enhanced reasoning in LLMs (Guo et al., 2025a; Team et al., 2025; OpenAI, 2025). In our approach, we construct preference pairs and assign verifiable rewards based on the correctness of the model’s judgments. By optimizing over the generated thinking traces, J1 encourages LLMs to spend more test-time compute before deriving scores and judgments.
@@ 6 Conclusion
We proposed J1, an RL recipe for training Thinking-LLM-as-a-Judge models. Our key innovation lies in converting the judgment task into a verifiable one for all kinds of task prompts, both verifiable and non-verifiable, and then optimizing the thoughts and judgments using an online RL method. We trained J1 at 8B, 32B, and 70B scales, exploring pointwise, pairwise, and multitask formulations. Our generalist J1 judge models outperformed all baselines at their respective sizes, with J1-Qwen-32B also surpassing the much larger R1 and o3 models on some benchmarks. Furthermore, we find that with only pairwise supervision, we can train pointwise judges that effectively mitigate position bias. Finally, we proposed a multitask training strategy that combines the strengths of pairwise judging (comparative context) and pointwise judging (position consistency), achieving the best overall performance and highlighting the potential Thinking-LLM-as-a-Judge.
Reproducibility statement
All J1 models are built upon open-weight Llama and Qwen models. The prompts used for data generation and training of J1 models are provided in the Appendix. Furthermore, the training code for J1 is based on the open-source Verl repository, and detailed hyperparameters to facilitate experiment reproducibility are listed in Appendix C.
Appendix A J1’s Thinking Traces
Here in Figure 7 and Figure 8 we show additional representative examples of J1’s thinking traces. The first is for verifiable math prompts and the third one is for a non-verifiable prompt.
Example 2 of J1’s Thinking Trace for a Verifiable Math Question <think>
Okay, let’s see. The user is asking about the area of an isosceles right triangle with hypotenuse endpoints at A(5,-5) and B(-1,-1). Both assistants provided answers, but I need to figure out which one is correct and better explained.
....So Assistant A’s answer is correct.
[…]
Therefore, Assistant A’s answer is correct with 13, while Assistant B’s is incorrect. The key error in B’s solution is the incorrect simplification of the product of the hypotenuse and the leg as 52 instead of the correct value. Assistant A’s approach is more straightforward and correct. Both solutions are step-by-step, but B’s mistake in the final calculation leads to an incorrect answer. So Assistant A’s response is better.
</think>
Figure 7: Thinking trace of J1 for a verifiable math prompt. The model analyzes Assistant B’s answer in a step-by-step manner, finds an error in the simplification step, and provides detailed feedback (highlighted in blue).
Example 3 of J1’s Thinking Trace for a Non-Verifiable Writing Prompt <think>
Okay, let’s see. The user is asking about the requirements for arc flash PPE, specifically mentioning maintenance, use, inspections, and shock protection areas like limited and restricted approach boundaries. They want the answer to be for only qualified persons.
Looking at Assistant A’s answer: They talk about qualified persons needing to identify PPE, ensure maintenance and inspections, understand shock protection areas, and train others. That covers the points mentioned. But wait, the user’s question is about the requirements for arc flash PPE, so maybe Assistant A is adding more about training others, which might be relevant but not directly the requirements for the PPE itself.
Assistant B’s answer starts by stating only qualified persons can work on exposed parts. They mention being trained, knowing equipment, hazards, PPE, inspections, and understanding the approach boundaries. This seems more focused on the qualifications and responsibilities of the person, which aligns with the user’s instruction to make the answer for only qualified persons.
The user’s question includes the requirements for PPE and shock protection areas. Both answers mention these, but Assistant B’s answer is more concise and directly addresses the requirements without adding extra info about training others. Also, Assistant B starts by stating that only qualified persons are allowed, which directly answers the user’s instruction. Assistant A’s answer might be more detailed but includes elements not directly asked for. So Assistant B’s answer is more precise and better follows the user’s specific request.
</think>
Figure 8: Thinking trace of J1 for a non-verifiable prompt.
Appendix B Prompt Templates
@@
Prompt for Creating Preference Pairs for J1 Training Below is a conversation between an user and an AI Assistant.
  {instruction}
The start of Assistant’s Answer
  {baseline response}
Please first generate a modified instruction that is highly relevant but not semantically identical to the instruction above from the user. Then write a high-quality answer which is a good response to the modified instruction but not a good response to the original user question. IMPORTANT: Please strictly follow the following format:
User Question Modified
<provide a modified instruction here>
The start of Assistant’s answer to the modified instruction
<provide a high-quality response to the modified instruction>
The end of Assistant’s answer to the modified instruction
Figure 14: Given a prompt and a baseline response, the prompt asks the LLM to generate a noisy instruction and then a response to that noisy instruction. The response to the noisy instruction subsequently is selected as the rejected response to the original instruction.
Appendix C Experimental Setup
@@
For training, the policy actor generates 5 rollouts per prompt using ancestral sampling with temperature $1.0$. Training regime uses a learning rate of 1e61e-6 (decayed to 3e73e-7 in later steps for pairwise J1-Llama-70B), and a total batch size of 512. The maximum sequence length is set to 4096 tokens for both inputs and outputs.
@@
We experimented with different KL coefficients from $\{0.1,0.01,0.001,0\}$ for J1-Llama-8B, and selected $0.01$ as the best-performing value based on development set accuracy. For J1-Llama-70B, we set the KL coefficient to 0 to encourage more exploration. Preliminarily experiment with entropy bonus during training showed degraded performance: the model tends to generate longer but more repetitive output. See Table 10 in Appendix D for comparison of KL penalty and entropy bonus.
[…]
Appendix D Additional Results
@@
Comparison of Different Pairwise-J1 Models. In Table 7, we compare the three Pairwise-J1 variants, that as part of the final answer, generate either: (i) only the final verdict, (ii) only real-valued scores for both responses, or (iii) both. We observe that predicting the verdict (without the scores) performs as well as other variants. Having access to scores, however, has other advantages e.g., in quantifying the degree of preference or to rank across multiple responses. Next, in Table 8, we compare the effect of training Pairwise-J1 models with both orders of data and with verdict consistency rewards. We observe that adding a consistency reward improves position-consistency accuracy of J1 models and consequently, also reduces the fraction of verdict flips.
Table 5: An example of a synthetically generated training pair for J1 training. The rejected response is a good response to the noisy instruction, thus making it a bad response to the original instruction.
Original Question Noisy Question Chosen Response Rejected Response
There are 3 objects arranged in a line. Someone says "the table is to the right of the bag, the table is to the left of the box and the bag is to the right of the box." Is the situation just described possible or impossible? Before answering, take a deep breath. Think step by step before giving a final answer. My survival depends on your answer being correct, so concentrate hard.Please analyze this new situation and determine whether it is possible or impossible. There are 3 objects arranged in a line. Someone says "the table is to the right of the bag, the box is to the left of the table and the bag is to the left of the box." Please analyze this new situation and determine whether it is possible or impossible. Let’s break down the statements: "The table is to the right of the bag." "The table is to the left of the box." "The bag is to the right of the box." From statement 3, we know the order is: box -> bag From statement 1, we know the order is: bag -> table Combining these two, we get: box -> bag -> table Now, let’s look at statement 2: "The table is to the left of the box." This statement contradicts our previous conclusion. If the table is to the left of the box, then the correct order should be: table -> box -> bag However, this contradicts statement 1, which says the table is to the right of the bag. Therefore, I conclude that the situation described is IMPOSSIBLE. The statements are mutually contradictory, and it’s not possible for all three statements to be true at the same time. " Let’s break down the given statements: The table is to the right of the bag. The box is to the left of the table. The bag is to the left of the box. From statement 3, we know the order is: bag -> box From statement 2, we know the order is: box -> table Now, let’s combine these orders: bag -> box -> table (from statements 2 and 3) Statement 1 says the table is to the right of the bag, which is consistent with this order. Conclusion Based on the analysis, it is POSSIBLE for all three statements to be true at the same time. The correct order is: bag -> box -> table.
Table 6: Results on RewardBench comparing our J1-Qwen-32B-MultiTask model with other top performing generative reward models. Results for most of these models are from the RewardBench leaderboard.
Models Overall Chat Chat-Hard Safety Reasoning
@@
CompassJudger CJ-1-32B (Cao et al., 2024) 85.4 97.8 65.6 86.1 92.2
facebook/Self-taught-evaluator-llama3.1-70B (Wang et al., 2024d) 90.0 96.9 85.1 89.6 88.4
Salesforce/SFR-nemo-12B-Judge-r (Wang et al., 2024b) 90.3 97.2 82.2 86.5 95.1
SF-Foundation/TextEval-OffsetBias-12B 91.0 91.9 86.6 92.0 93.6
Reward Reasoning Model (Guo et al., 2025b) 91.2 94.7 81.1 90.7 98.3
RM-R1 (Chen et al., 2025b) 91.4 95.3 83.1 91.9 95.2
AtlaAI/Selene-1 (Alexandru et al., 2025) 92.4 97.8 84.0 92.2 95.7
R-I-S-E/RISE-Judge-Qwen2.5-32B (Yu et al., 2025a) 92.7 96.6 83.3 91.9 98.8
Salesforce/SFR-LLaMa-3.1-70B-Judge-r (Wang et al., 2024b) 92.7 96.9 84.8 91.6 97.6
Skywork/Skywork-Critic-Llama-3.1-70B (Shiwen et al., 2024) 93.3 96.6 87.9 93.1 95.5
SF-Foundation/TextEval-Llama3.1-70B 93.5 94.1 90.1 93.2 96.4
J1-Qwen-32B-MultiTask (Ours) 93.6 96.4 89.5 90.5 98.1
Pairwise-J1 8B Variants Overall PPE RewardBench JudgeBench RM-Bench FollowBenchEval
w/ Verdict : $(x,a,b)\rightarrow(t,y)$ 63.9 59.8 85.7 42.0 73.4 48.3
w/ Scores: $(x,a,b)\rightarrow(t,s_{a},s_{b})$ 63.4 60.2 85.8 41.7 72.3 46.3
w/ Scores&Verdict: $(x,a,b)\rightarrow(t,s_{a},s_{b},y)$ 61.7 59.5 85.1 41.4 71.5 41.0
Comparison of Different Decoding Hyperparameters for J1 Models. In Table 9, we compare greedy decoding to temperature sampling of J1 models. We find that our models are robust to such choices, exhibiting consistent performance with negligible variance.
Comparison of J1 with a Scalar RM Trained on Same Data. In Table 11, we compare J1 to a scalar RM trained on the same base model (Llama-3.1-8B-Instruct) using the same training data. We observe that J1 outperforms the corresponding scalar model on four out of five benchmarks with the maximum improvement coming in the hardest RM-Bench benchmark.
Table 8: Results on PPE Correctness and JudgeBench comparing different position-bias mitigation strategies for Pairwise-J1. Consistent Accuracy: examples are judged correctly in both response orders. Verdict Flip: cases where the pairwise verdict changes when response order is swapped.
Pairwise-J1 8B Variants PPE Correctness JudgeBench
(a, b) (b, a) Consistent Verdict- (a, b) (b, a) Consistent Verdict-
Acc $\uparrow$ Acc $\uparrow$ Acc $\uparrow$ Flip $\downarrow$ Acc $\uparrow$ Acc $\uparrow$ Acc $\uparrow$ Flip/Ties $\downarrow$
Llama-3.1-8B-Instruct 54.7 54.1 30.2 44.1 67.4 42.3 32.3 37.4
Random Single-order Data 58.3 57.6 38.3 36.7 48.3 59.4 36.6 32.9
Both-order data 59.2 58.4 39.1 36.8 63.1 51.4 42.0 27.7
Verdict Consistency Reward 58.4 58.2 43.9 28.7 52.3 64.6 45.4 26.0
Table 9: Comparison of Greedy Decoding and Temperature Sampling showing the robustness of J1 models to different decoding temperatures.
Model Decoding Temperature PPE Correctness
J1-Llama-70B Greedy (with t=0.0) 72.9
J1-Llama-70B Temperature Sampling (8 seeds with t=0.6) 72.8 $\pm$ 0.1
J1-Qwen-32B-MultiTask Greedy (with t=0.0) 76.8
J1-Qwen-32B-MultiTask Temperature Sampling (8 seeds with t=0.6) 77.0 $\pm$ 0.1
Pairwise-J1 8B Variants Overall PPE RewardBench JudgeBench RM-Bench FollowBenchEval
w/ KL Penalty 63.9 59.8 85.7 42.0 73.4 48.3
w/o KL Penalty 61.1 59.6 84.9 42.9 69.8 48.3
w/o Entropy Bonus 63.9 59.8 85.7 42.0 73.4 48.3
w/ Entropy Bonus 58.3 59.1 84.1 39.4 71.7 42.4
Table 11: Comparison of J1 with a scalar RM trained on same data.
Model Overall PPE RewardBench JudgeBench RM-Bench FollowBenchEval
Bradley-Terry (scalar) 66.6 57.5 82.6 74.0 51.1 67.8
J1-Llama-8B (generative) 69.2 59.8 85.7 73.4 58.5 68.8
Table 12: Test-time scaling of Pointwise-J1 models on PPE Correctness. Judgments are made based on the average scores of the responses.
Pointwise-J1 Overall MMLU-Pro MATH GPQA MBPP-Plus IFEval
J1-Llama-70B - Greedy 65.0 73.4 77.4 58.9 60.7 54.6
J1-Llama-70B - Sampling 64.9 74.0 79.7 58.2 55.7 59.7
J1-Llama-70B (Mean-Score@32) 74.8 81.2 87.6 67.3 81.9 70.8
Table 13: Test-time scaling of Pairwise-J1 models on PPE Correctness. Judgments are made based on majority vote over multiple verdicts.
Pairwise-J1 Overall MMLU-Pro MATH GPQA MBPP-Plus IFEval
J1-Llama-70B - Greedy 72.9 79.0 86.0 65.9 66.0 67.3
J1-Llama-70B (SC@32) 73.7 79.9 88.1 66.5 66.5 67.2